OpenStudy (anonymous):

Find the largest δ such that if 0 < |x – 1| < δ then |f(x) – 1| < 0.1?

OpenStudy (anonymous):

\[f(x) = 2- \frac{ 1 }{ x }\]

OpenStudy (amistre64):

what have you worked out so far?

OpenStudy (anonymous):

@amistre64 I did f(x)=2-1/x absolute value of 2-1/x < 0.1 0< absolute value of x-1< delta 1- 1/x < 0.1 Here's where I got stuck. Di I do it wrong?

OpenStudy (michele_laino):

we have to solve these two inequalities: \[\Large \left| {1 - \frac{1}{x}} \right| < \frac{1}{{10}}\]

OpenStudy (amistre64):

your starting fine, the idea is to convert the range into the domain, which is just algebra worked over ... \[|f(x)-1|<.1\] \[-.1<f(x)-1<.1\] \[.9<f(x)<1.1\] \[.9<2 - \frac 1x<1.1\] \[.9-2<- \frac 1x<1.1-2\] \[2-.9> \frac 1x>2-1.1\] \[\frac1{2-.9}< x<\frac1{2-1.1}\] \[\frac1{2-.9}-1< x-1<\frac1{2-1.1}-1\]

OpenStudy (amistre64):

we have 2 deltas to try .... test them out to see what fits.

OpenStudy (amistre64):



OpenStudy (amistre64):

the smaller value is the safest bet to me ...

OpenStudy (anonymous):

I've gotten -0.09 <x-1<0.11

OpenStudy (amistre64):

-1/11 < x < 1/9

OpenStudy (anonymous):

o sorry

OpenStudy (amistre64):

\[2-\frac1{1-\frac1{11}}=1-.1\] \[2-\frac1{1+\frac1{9}}=1+.1\]

OpenStudy (amistre64):

but since 1/9 is greater than 1/11 we cant use x-1/9 and still be in our range of .9 to 1.1

OpenStudy (amistre64):

\[2-\frac1{1-\frac1{9}}=1-.125\] we are out of the bounds of our limitation

OpenStudy (anonymous):

That'd give us 0.875 or 7/8

OpenStudy (amistre64):

correct, which is smaller than .9 we need to remain within (.9, 1.1) .875 < (.9, 1.1)

OpenStudy (anonymous):

right

OpenStudy (amistre64):

if we use the delta = 1/11 then the function fits in both cases. so its the largest value we can use.

OpenStudy (anonymous):

which gives us 0.0909....

OpenStudy (amistre64):

yes if they want a decimal representation of it

OpenStudy (anonymous):

I checked it in my assignment online for that question, and 1/11 is the correct answer.

OpenStudy (anonymous):

They take integers, fractions or decimals

OpenStudy (amistre64):

well of course it is .. :)

OpenStudy (anonymous):

@amistre64 thank you for the walk through. I just got stuck whiles working it out :)

OpenStudy (amistre64):

youre welcome, ita just an exercise in algebra, converting range to the domain and picking the smaller result

OpenStudy (anonymous):

u sir just got yourself a medal ;)

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank u